A can is dropped from rest at a height of 40.0 m above ground. How much time does it spend in the bottom half of its fall?


  1. 0.873 s
  2. 0.783 s
  3. 0.837 s
  4. 0.738 s
1
Expert's answer
2022-01-23T16:15:00-0500

"t_0=\sqrt\frac{2s}g=2.857~s,"

"t_1=\sqrt\frac sg=2.02~s,"

"t_2=t_0-t_1=0.837~s."